OpenStudy (anonymous):

If 45g of LiF dissolved in 1.8kg of water, what would be the expected change in boiling point? The boiling point constant for water (Kb) is 0.51 degrees C/m. - 0.49 degrees C - 0.98 degrees C - 1.9 degrees C - 3.5 degrees C Im not sure which one it is I got both 0.49 and 0.98 Please Help. I will medal.

OpenStudy (aaronq):

What calculations did you perform?

OpenStudy (anonymous):

I ended up going with the first answer 0.49 based on my calculations

OpenStudy (aaronq):

did you have \(\large \Delta T=i*K_b*m=2*(0.51~^oC/m)*\dfrac{(\frac{45~g}{25.9~g/mol})}{1.8~kg}\)

OpenStudy (anonymous):

No and when I solve yours I get 25 which isn't a choice

OpenStudy (aaronq):

you're solving it incorrectly, because i get 0.984555984555984583356

OpenStudy (anonymous):

I'm so lost

OpenStudy (anonymous):

I don't know who's right

OpenStudy (aaronq):

I'm right lol you got 0.49 because you didn't take into account that i=2

OpenStudy (anonymous):

Okay thank you so much

OpenStudy (anonymous):

CaN I ask another

OpenStudy (aaronq):

no problem, and sure, shoot.

OpenStudy (anonymous):

A 65 gram of some unknown metal at 100.0 deg C is added to 100.8grams at 22 degrees Celsius. The temperature of the water rises 27.0 degrees C. Of the specific heat capacity of liquid water is 4.18 J/degC*g, what is the specific heat of the metal?

OpenStudy (anonymous):

The choices are -2.25 J/(C*g) -1.75 J/(C*g) -0.444 J/(C*g) -0.324 J/(C*g)

OpenStudy (anonymous):

I don't know Where to start @arronq and how to get the answer

OpenStudy (aaronq):

So this is a little complicated but bare with me. First you use the mass of the water, the specific heat capacity and the change of temperature the water had to find the heat lost by the metal. Now that you have the heat lost by the metal, you can use the mass of the metal and the temperature change of the metal (along with the heat you found in the first part) to solve for the specific heat capacity. The equation is \(\sf q=m*C_p*\Delta T\) q= heat m=mass Cp=specific heat capacity \(\Delta T\)=change in Temp

OpenStudy (anonymous):

Based on what your saying and what I do know I'm getting 0.444 with my calculations. Is that correct?

OpenStudy (aaronq):

show your calculations

OpenStudy (anonymous):

I can't I don't have a camera and my computers to old to type them in

OpenStudy (aaronq):

so how am i supposed to see if you're doing it right?

OpenStudy (anonymous):

I think I am I just need to double check the final response with you and if I get it wrong I'll try it a different way until I'm doing it right

OpenStudy (aaronq):

its right

OpenStudy (anonymous):

Okay thank you so much

OpenStudy (aaronq):

no problem